OpenStudy (anonymous):

using implicit differentiation to find y' for 2y2-3x3-5=0 and evaluate at (x,y)=(1,2)

OpenStudy (anonymous):

2y^2-3x^3-5=0 2ydy/dx-9x^2=0 2ydy/dx=9x^2 dy/dx=9x^2/2y at (x,y)=(1,2) y'=9^2/2(2) =81/4

OpenStudy (anonymous):

4yy'-9x^2=0 y'=9x^2/4y @ (1,2) y' =9/8

OpenStudy (anonymous):

thank you